OpenStudy (anonymous):

compute the difference quotient of 2x^2+4x-1 a)4x+4 b)2x+2h+2 c)4x+2h+4 d)4x-2h-4

OpenStudy (anonymous):

\[\frac{\left(2(x+h)^2+4(x+h)-1\right)-\left(2x^2+4x-1\right)}{h}\\ \frac{2x^2+4xh+2h^2+4x+4h-1-2x^2-4x+1}{h}\\ \frac{4xh+2h^2+4h}{h}\\ 4x+2h+4 \]